OpenStudy (anonymous):

Solve 2/3 (c-18) = 7

OpenStudy (anonymous):

do you mean (2/3) (c-18) = 7 or 2 / ( 3 (c-18) ) = 7?

OpenStudy (anonymous):

In other words, is (c-18) a standalone term multiplied by 2/3, or is it in the denominator of 2/3?

OpenStudy (anonymous):

c-18 stand alone

OpenStudy (anonymous):

Assuming the first case, you would have (2/3) (c-18) = 7 (2(c-18) ) / (3) = 7 (2c-36) / 3 = 7 2c-36 = 7 * 3 2c-36 = 21 2c = 21+36 2c = 57 c = 57 / 2 c = 28.5

OpenStudy (anonymous):

c-18 = 7 x 3/2 = 21/2 c= 18 +21/2 = 28.5